OpenStudy (hba):

Show that for a projectile d^2(v^2)/dt^2=2g^2

OpenStudy (aravindg):

do you know general equation of a projectile?

OpenStudy (hba):

I'll derive it if you want :(

OpenStudy (callisto):

Ok, you need to derive it. Derive v^2

OpenStudy (hba):

@Callisto Range*g * sin 2 θ = V^2

OpenStudy (callisto):

range = ?

OpenStudy (hba):

okay what do you want now :/

OpenStudy (hba):

If i knew everything i wouldn't have posted the ques.

OpenStudy (callisto):

You can derive the range too... But... My approach was to find \(v_x\) and \(v_y\) So, \(v^2 = v_x +v_y\)

OpenStudy (callisto):

\(v_x\) and \(v_y\) will be in terms of u, g and t I think you can take u and g as constants.

OpenStudy (callisto):

u = initial velocity

OpenStudy (hba):

V2 =V; +V;. For a projectile Vx is constant, so we need only evaluate ~(V;)/dt2. The first derivative is 2vy dvy / dt = -2vyg. The derivative of this ( the second derivative) is -2g dvy / dt = 2g2.

OpenStudy (hba):

V^2=Vx^2+Vy^2

OpenStudy (callisto):

You should explain why dvy / dt = g then.

OpenStudy (hba):

Okay so,